# Comparing Machine Learning Algorithms by Union-Free Generic Depth

### 20 Dec 2023 v1

@@ 4.1 Properties of $\mathscr{S}$
We begin with introducing some properties of $\mathscr{S}$. Later, these properties are used to analyze the (empirical) ufg depth and to improve the computation. First of all, we want to introduce a second equivalent definition of Condition (C1) and Condition (C2). This will later help to prove the connectedness property, see Theorem 3, and to improve the implementation, see Section 5.
@@
Based on Lemma 1 and the definition of the closure operator on posets, see Definition 1, we can immediately obtain a description of the ufg elements w.r.t. some set $S\in\mathscr{S}$. Definition 5 states that $q$ is an ufg element w.r.t. $S\in\mathscr{S}$ if and only if each poset $p\in S$ contributes to $q$ in a manner which only this poset $p$ can achieve. More precisely, for each $p\in S$ we have that either $p$ has edge $(y_{1},y_{2})$ which only $p$ and $q$ share or $p$ is the only poset in $S$ which does not have edge $(y_{1},y_{2})$ which also $q$ does not have. Note that if we ensured that each poset $p\in S$ contributes to a poset $q\in\gamma(S)$ then this $q$ is an ufg element w.r.t. $S$ and $S$ fulfills Condition (C1) and (C2). Based on this discussion, we define the following two sets:
@@
Using these definitions we obtain by Lemma 1 and its proceeding discussion the following corollary.
@@
After discussing the different definitions of sets in $\mathscr{S}$, let us now consider the claims made in Section 3. These claims are that sets of cardinality one cannot be union-free and generic, and that only special cases of sets of cardinality two are union-free and generic. These two claims are proved in the following theorem. Furthermore, this theorem shows that the sets in $\mathcal{S}$ are connected in the sense that for every $S\in\mathscr{S}$ with #$\#S=m\geq 3$, there exists a $S_{m}\in\mathscr{S}$ such that $S_{m}\subsetneq S$ and #$\#S_{m}=m-1$.
@@
Theorem 3 Part 1 gives us directly a lower bound for the cardinality of all sets $S\in\mathscr{S}$. For the upper bound, we use a complexity measure of $\mathscr{S}$, the Vapnik-Chervonenkis dimension (VC dimension for short), see Vapnik and Chervonenkis (2015). The VC dimension of a family of sets $\mathscr{C}$ is the largest cardinality of a set $A$, such that $A$ can still be shattered into the power set of $A$ by $\mathscr{C}$ [^15] To be more precise: The intersection between a set $A$ and a family of sets $\mathscr{C}$ is defined by $A\cap\mathcal{C}=\{A\cap C\mid C\in\mathscr{C}\}$. We say that a set $A$ can be shattered (by $\mathscr{C}$) if #$\#(A\cap\mathscr{C})=2^{\#A}$ holds. The VC dimension of $\mathscr{C}$ is now defined as $vc=\max\{\#A\mid(A\cap\mathscr{C})=2^{\#A}\}.$. With this, we obtain an upper bound for all $S\in\mathscr{S}$ is given by #${\#S\leq vc}$, with $vc$ the VC dimension of the closure system $\gamma(2^{\mathcal{P}})$. Note that in our case of posets, the VC dimension is small compared to the number of all posets.
@@
To prove #$\#S\leq vc$ take an arbitrary subset $Q=\{p_{1},\ldots,p_{k}\}\in\mathscr{S}$ of size $k>vc$. Then this subset is not shatterable and thus there exists a subset $R\subseteq Q$ that cannot be obtained as an intersection of $Q$ and some $\gamma(A)$ with $A\subseteq\mathcal{P}$. In particular, this holds for $R=A$. Thus, $R\neq\gamma(R)\cap Q$ and with the extensitivity of $\gamma$ we get $R\subsetneq\gamma(R)\cap Q$. This means that there exists an order $p_{i}$ in $\gamma(R)\cap Q\backslash R$ for which the formal implication $R\rightarrow\{p_{i}\}$ holds. Thus, (because of the Armstrong rules, cf., (Armstrong, 1974, p. 581)) the order $p_{i}$ is redundant in the sense of $Q\backslash\{p_{i}\}\rightarrow Q$ and thus $Q$ is not minimal with respect to $\gamma$. Therefore, $Q$ is not in $\mathscr{S}$ which completes the proof.
@@
In concrete applications, one usually does not observe all possible posets. Therefore, many summands in the definition of the ufg depth (see Definition 4) are zero. Thus, one can restrict the analysis to $\mathscr{S}_{obs}:=\{S\in\mathscr{S}\mid\mbox{ all }p\in S\mbox{ are observed}\}.$ Then a similar argumentation shows that #$\#S\leq vc_{obs}$ where $vc_{obs}$ is the VC dimension of $\gamma(2^{\mathcal{P}_{obs}})\cap\mathcal{P}_{obs}$ and $\mathcal{P}_{obs}$ is the set of all observed posets.
@@ 4.2 Properties of the (Empirical) UFG Depth
Proof 6
We simply give two data sets $\mathcal{D}=(p_{1},p_{2},p_{3})$ and $\tilde{\mathcal{D}}=(\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3})$ on the basic set $M=\{y_{1},y_{2},y_{3}\}$ with the same sum-statistics but different associated depth functions: Let $p_{1},p_{2}$ and $p_{3}$ be given as the transitive reflexive closures of $\{(y_{1},y_{2})\}$; $\{(y_{1},y_{2}),(y_{1},y_{3})\}$ and $\{(y_{2},y_{3}),(y_{1},y_{3})\}$, respectively. Let $\tilde{p}_{1}$, $\tilde{p}_{2}$ and $\tilde{p}_{3}$ be the transitive reflexive closure of $\{(y_{1},y_{2})\}$; $\{(y_{1},y_{3})\}$ and $\{(y_{1},y_{2}),(y_{2},y_{3})\}$, respectively. Then both data sets have the same sum-statistics $w_{(y_{1},y_{2})}=w_{(y_{1},y_{3})}=2$; $w_{(y_{1},y_{3})}=1$ and $w_{(y_{i},y_{j})}=0$ for all other $y_{i}\neq y_{j}$. But the ufg depth of $p_{1}=\tilde{p}_{1}$ is $1/2$ w.r.t. the first data set but $7/10$ w.r.t the second data set. The corresponding code can be found at the link mentioned in Footnote $1$.
@@
Proof 8
First, note that for $S\in\mathscr{S}$, where there exists an $\tilde{p}\in S$ such that $\nu_{n}(\tilde{p})=0$, $S$ contributes nothing to $D_{n}$. So one can replace $\mathscr{S}$ in the definition of $D_{n}$ by $\tilde{\mathscr{S}}=\{S\cap\{p_{1},\ldots,p_{n}\}\mid S\in\mathscr{S}\}$. The reduced set $\tilde{\mathscr{S}}$ is used to show the claims.
The proof of Claim 1., 2., 3. and 4. are analogous. Hence, here we provide only the proof of Claim 1. Let $(y_{1},y_{2})\in M\times M$ such that for all $i\in\{1,\ldots,n\}$ $(y_{1},y_{2})\in p_{i}$ and let $p\in\mathcal{P}$ such that $(y_{1},y_{2})\not\in p$. Let $S\in\mathscr{S}\cap\{p_{1},\ldots,p_{n}\}$ and take a closer look at Equation-part (3) of Lemma 5. Since $(y_{1},y_{2})\not\in p$, $S$ cannot be an element of the intersection of (3). Thus, $\{S\cap\{p_{1},\ldots,p_{n}\}\in\mathscr{S}\mid p\in\gamma(S)\}$ is empty and with the comment above we get that $D_{n}(p)=0$.
The last properties have summarized how the structure of a sample is reflected in the ufg depth and vice versa. Finally, we take a look at the consistency of the empirical ufg depth $D_{n}$. This means that $D_{n}$ converges uniformly to $D$ almost surely under the assumption of observing iid samples, see Theorem 9.
@@
Let for all $n\in\mathbb{N}$ the sample $(p_{1},\ldots,p_{n})$ be iid according to distribution $\nu\in\mathcal{M}$. Then the corresponding empirical depth function $D_{n}$ almost surely uniformly to $D(\cdot,\nu)$ for $n$ going to infinity.
Proof 9
Due to the iid assumption and the law of large numbers, we know that for every $p\in\mathcal{P},\>\left\lVert\nu_{n}(p)-\nu(p)\right\rVert\overset{n\to\infty}{\to}0$ almost surely (a.s). Since #$\#\mathcal{P}$ is finite, we get that $\nu_{n}$ also converges a.s. uniformly to $\nu$. Finally, we use that $D_{n}$ and $D$ are both the same finite composition of $\nu_{n}$ and $\nu$, respectively, and we obtain $\sup_{p\in\mathcal{P}}\left\lVert D_{n}(p)-D(p)\right\rVert\overset{n\to\infty}{\to}0\text{ almost surely.}$.
@@ 5 Implementation
In this section, we discuss the difficulties and corresponding solution approaches in computing $D_{n}$. Therefore, let $\underline{p}=(p_{1},\ldots,p_{n})$ be a sample of posets. The naive approach is to just check all subsets of $\{p_{1},\ldots,p_{n}\}$ whether they are in $\mathscr{S}_{obs}:=\{S\in\mathscr{S}\mid\mbox{ all }p\in S\mbox{ are observed}\}$ is very time-consuming, especially since the subsets that are elements of $\mathscr{S}_{obs}$ can be very sparse in $2^{\{p_{1},\ldots,p_{n}\}}$. Moreover, it is difficult to test whether a subset is an element of $\mathscr{S}_{obs}$ or if it is not an element by taking Conditions (C1) and (C2) as a basis. Finally, even if we are able to compute the entire family $\mathscr{S}_{obs}$, computing all possible poset $p$ on a set $M$ to obtain some insights on $D_{n}$ is for even small $M$ with #$\#M\geq 8$ very computation intensive and for larger #$\#M$ it is currently not possible. [^16] Even computing the entire number of possible posets is not computable for #$\#M>18$, see https://oeis.org/A001035 (Accessed 10.11.2023). We address all these issues in the next paragraphs.
@@
where $\frac{w_{S}}{c_{n}}$ for $S\in\mathscr{S}_{obs}$ are the weights given by duplicates and the cardinality of $S$ together with the normalization constant, see Definition 4, and $x_{i}\in\{0,1\}$ for $i\in\mathscr{S}_{obs}\cup(M\times M)$. The first #$\#\mathscr{S}_{obs}$ elements of $x$ describe the union-free generic set, and the next #${\#M\cdot\#M}$ elements guarantee that we only discuss posets. First, we need to ensure that only relations that define a poset are discussed. Therefore one part of the constraint matrix represents only the reflexive, transitive, and antisymmetric constraints on $x_{(a,b)}$ with $(a,b)\in M\times M$. So in this part, the matrix entries of the first #$\#\mathscr{S}_{obs}$ columns are all zero. Second, we have to include in the binary linear program that a set $S\in\mathscr{S}_{obs}$ can only be included in the maximum if the considered poset lies in the intersection of all posets of $S$. This can be ensured by a row in the constraint matrix that imposes a structure on $x_{(a,b)}$ with $(a,b)\in M\times M$ when this set $S$ is included in the maximization of the objective function. Similarly, we ensure that a set $S\in\mathscr{S}_{obs}$ is only counted if the poset considered is a subset of the union of all posets in $S$. Now, computing the maximum leads to the highest depth value and a poset that has that highest depth value. However, if we are interested in the smallest depth value, this binary linear program does not work, because we have not yet enforced that a set $S\in\mathscr{S}_{obs}$ which has the currently analyzed poset in its closure must be represented in the calculation of the objective function. Therefore, we need to include further constraints so that when a poset is used to minimize the objective, every set $S\in\mathscr{S}_{obs}$ that has that poset in its closure is counted in the objective sum.
@@
All in all, we improved the computation compared to the naive approach by using the knowledge provided in Section 4. By the two bounds on the cardinality of the sets in $\mathscr{S}_{obs}$ we get an idea of the worst and best case of the computation time. By further using Lemma 1 with Corollary 2 together the connectedness property, see Theorem 3 Part 3, we could decrease the computation time. Although we currently cannot calculate the exact amount of this reduction in general as this depends on the complexity of the data set used. Note that even the upper bound is not fixed, but depends on the structure of the data set. Finally, by stating a binary linear program, we achieved that even if not all posets are computable, we get the highest and lowest ufg depth values with representative posets. A summarizes the computation time and complexity of the two examples in the next section.

### 21 Feb 2024 v3

@@ 5.1 Properties of $\mathscr{S}$
We begin with introducing some properties of $\mathscr{S}$. Later, these properties are used to analyze the (empirical) ufg depth and to improve the computation. First, we want to introduce a second equivalent definition of Condition (C1) and Condition (C2). This says that $S\in\mathscr{S}$ if and only if there exists a poset $q$ such that every poset $p\in S$ contributes to $q$. More precisely, for every $p\in S$ we have that either $p$ has an edge $(y_{1},y_{2})$ which only $p$ and $q$ share, or $p$ is the only poset in $S$ which does not have an edge $(y_{1},y_{2})$ which also $q$ does not have. Note that if we have ensured that every poset $p\in S$ contributes to a poset $q\in\gamma(S)$, then $S$ satisfies Conditions (C1) and (C2). This will help later to prove the connectedness property, see Theorem 3, give an upper bound for #$\#S$ with $S\in\mathscr{S}$, see Theorem 5, and to improve the implementation, see Section 6.
@@
As we pointed out above, if $q$ is an ufg element w.r.t. some set $S\in\mathscr{S}$, then each poset $p\in S$ has two different ways of contributing to $q$: Either it has an edge that only $p$ and $q$ share, or it has an edge that only $p$ and $q$ do not have. With this we immediately get a description of the ufg elements w.r.t. some set $S\in\mathscr{S}$. Therefore we define the following two sets, which sort the posets into two sets w.r.t. $S\in\mathscr{S}$. One consists of those posets that have an edge that all other posets in $S$ do not have, and conversely the other set contains all posets that do not have an edge that all other posets in $S$ have. Note that these two sets are not necessarily disjoint.
@@
Using these definitions we obtain by Lemma 1 and its corresponding discussion a concrete definition of an ufg element w.r.t. $S$ if $S\in\mathscr{S}$, see the next corollary.
@@
After discussing the different definitions of sets in $\mathscr{S}$, let us now consider the claims made in Section 4. These claims are that sets of cardinality one cannot be union-free and generic, and that only special cases of sets of cardinality two are union-free and generic. These two claims are proved in the following theorem. Furthermore, this theorem shows that the sets in $\mathcal{S}$ are connected in the sense that for every $S\in\mathscr{S}$ with #$\#S=m\geq 3$, there exists $S_{m}\in\mathscr{S}$ such that $S_{m}\subsetneq S$ and #$\#S_{m}=m-1$.
@@
Theorem 3 Part 1 gives us directly a lower bound for the cardinality of all sets $S\in\mathscr{S}$. For the upper bound, we use a complexity measure of $\mathscr{S}$, the Vapnik-Chervonenkis dimension (VC dimension for short), see Vapnik and Chervonenkis (2015). The VC dimension of a family of sets $\mathscr{C}$ is the largest cardinality of a set $A$, such that $A$ can still be shattered into the power set of $A$ by $\mathscr{C}$ [^15] To be more precise: The intersection between a set $A$ and a family of sets $\mathscr{C}$ is defined by $A\cap\mathcal{C}=\{A\cap C\mid C\in\mathscr{C}\}$. We say that a set $A$ can be shattered (by $\mathscr{C}$) if #$\#(A\cap\mathscr{C})=2^{\#A}$ holds. The VC dimension of $\mathscr{C}$ is now defined as $vc=\max\{\#A\mid(A\cap\mathscr{C})=2^{A}\}.$. With this, we obtain an upper bound for all $S\in\mathscr{S}$ is given by #${\#S\leq vc}$, with $vc$ the VC dimension of the closure system $\gamma(2^{\mathcal{P}})$. Note that in our case of posets, the VC dimension is small compared to the number of all posets.
@@
To prove #$\#S\leq vc$ take an arbitrary subset $Q=\{p_{1},\ldots,p_{k}\}\in\mathscr{S}$ of size $k>vc$. Then this subset is not shatterable because $vc$ is the largest cardinality of a shatterabel set. Thus there exists a subset $R\subseteq Q$ that cannot be obtained as an intersection of $Q$ and some $\gamma(A)$ with $A\subseteq\mathcal{P}$. In particular, this holds for $R=A$. Thus, $R\neq\gamma(R)\cap Q$ and with the extensitivity of $\gamma$ we get $R\subsetneq\gamma(R)\cap Q$. This means that there exists an order $p_{i}$ in $\gamma(R)\cap Q\backslash R$ for which the formal implication $R\rightarrow\{p_{i}\}$ holds. Thus, (because of the Armstrong rules, cf., (Armstrong, 1974, p. 581)) the order $p_{i}$ is redundant in the sense of $Q\backslash\{p_{i}\}\rightarrow Q$ and thus $Q$ is not minimal with respect to $\gamma$. Therefore, $Q$ is not in $\mathscr{S}$ which completes the proof.
@@
In concrete applications, one usually does not observe all possible posets. Therefore, many summands in the definition of the empirical ufg depth (see Definition 4) are zero. Thus, one can restrict the analysis to $\mathscr{S}_{obs}:=\{S\in\mathscr{S}\mid\mbox{ all }p\in S\mbox{ are observed}\}.$ Then a similar argumentation shows that #$\#S\leq vc_{obs}$ where $vc_{obs}$ is the VC dimension of $\gamma(2^{\mathcal{P}_{obs}})\cap\mathcal{P}_{obs}$ and $\mathcal{P}_{obs}$ is the set of all observed posets.
Beyond the bound for the size of a set $S\in\mathscr{S}$ that is given by the VC dimension and that is generally valid, we can additionally give another bound for the special case of poset data that is sharper than the VC bound in certain cases.
Theorem 5
Let $m:=\#M\geq 3$. Then, for every $S\in\mathscr{S}$ we have that #$\#S\leq m(m-1)/2$.
Proof 5
[…]
  • (i)
    contribute in $q$ by providing an edge $(x_{i},x_{j})\in q\cap p$ that no other poset in $S$ has, or
  • (ii)
    to contribute in $q$ by not having an edge $(x_{i},x_{j})\notin q$ and $(x_{i},x_{i})\notin p$ that every other poset has.
This means that $p$ has a unique existence-characteristic w.r.t. edge $(x_{i},x_{j})$ and $S$, while all other posets in $S$ agree to have the opposite existence-characteristic w.r.t. $(x_{i},x_{j})$ and $S$. That is, if $(x_{i},x_{j})\in p$, for all $\tilde{p}\in S\setminus\{p\}$ we have $(x_{i},x_{j})\not\in\tilde{p}$ and vice versa for $(x_{i},x_{j})\not\in p$.
Since by Corollary 2 each poset must somehow contribute uniquely to $q$, we get that each edge can be used by only one single poset in $S$. This follows from the fact that #$\#S\geq 3$ and all other posets $\tilde{p}\in S\setminus\{p\}$ need to agree on the opposite existence-characteristic w.r.t. $(x_{i},x_{j})$. With this we obtain that #$\#S\leq m(m-1)$ (since the reflexive part holds by default for every poset).
We continue with the above poset $p\in S$ and the corresponding edge $(x_{i},x_{j})$ that uniquely contributes to $q$ by $p$. We show that the inverse edge $(x_{j},x_{i})$ cannot be a contributing element for any $\tilde{p}\in S\setminus\{p\}$. To show this, we split the proof into two cases:
[…]
With these two cases we have that only half of all possible edges can be used and therefore we have shown that #$\#S\leq m(m-1)/2$.
Remark 3
The bound of Theorem 5 is tight in the sense that there exists a set $S\in\mathscr{S}$ with cardinality $m(m-1)/2$, namely the set
$S:=\{\{(x_{i},x_{j})\}\mid i,j\in\{1,\ldots,m\},i<j\}.$
@@ 5.2 Properties of the (Empirical) UFG Depth
Proof 7
We simply give two data sets $\mathcal{D}=(p_{1},p_{2},p_{3})$ and $\tilde{\mathcal{D}}=(\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3})$ on the basic set $M=\{y_{1},y_{2},y_{3}\}$ with the same sum-statistics but different associated depth functions: Let $p_{1},p_{2}$ and $p_{3}$ be given as the transitive reflexive closures of $\{(y_{1},y_{2})\}$; $\{(y_{1},y_{2}),(y_{1},y_{3})\}$ and $\{(y_{2},y_{3}),(y_{1},y_{3})\}$, respectively. Let $\tilde{p}_{1}$, $\tilde{p}_{2}$ and $\tilde{p}_{3}$ be the transitive reflexive closure of $\{(y_{1},y_{2})\}$; $\{(y_{1},y_{3})\}$ and $\{(y_{1},y_{2}),(y_{2},y_{3})\}$, respectively. Then both data sets have the same sum-statistics $w_{(y_{1},y_{2})}=w_{(y_{1},y_{3})}=2$; $w_{(y_{1},y_{3})}=1$ and $w_{(y_{i},y_{j})}=0$ for all other $y_{i}\neq y_{j}$. But the ufg depth of $p_{1}=\tilde{p}_{1}$ is $1/2$ w.r.t. the first data set but $7/10$ w.r.t the second data set. The corresponding code can be found at the link mentioned in Footnote 5.
@@
Proof 9
First, note that for $S\in\mathscr{S}$, where there exists an $\tilde{p}\in S$ such that $\nu_{n}(\tilde{p})=0$, $S$ contributes nothing to $D_{n}$. So one can replace $\mathscr{S}$ in the definition of $D_{n}$ by $\mathscr{S}_{obs}$, see Remark 2. The reduced set $\mathscr{S}_{obs}$ is used to show the claims.
The proof of Claim 1., 2., 3. and 4. are analogous. Hence, here we provide only the proof of Claim 1. Let $(y_{1},y_{2})\in M\times M$ such that for all $i\in\{1,\ldots,n\}$ $(y_{1},y_{2})\in p_{i}$ and let $p\in\mathcal{P}$ such that $(y_{1},y_{2})\not\in p$. Let $S\in\mathscr{S}_{obs}$ and take a closer look at Equation-part (3) of Lemma 6. Since $(y_{1},y_{2})\not\in p$, $S$ cannot be an element of the intersection of (3). Thus, $\{S\in\mathscr{S}_{obs}\mid p\in\gamma(S)\}$ is empty and with the comment above we get that $D_{n}(p)=0$.
The last properties have summarized how the structure of a sample is reflected in the ufg depth and vice versa. Finally, we take a look at the consistency of the empirical ufg depth $D_{n}$. This means that $D_{n}$ converges uniformly to $D$ almost surely under the assumption of observing i.i.d. samples, see Theorem 10.
@@
Let for all $n\in\mathbb{N}$ the sample $(p_{1},\ldots,p_{n})$ be i.i.d. according to distribution $\nu\in\mathcal{M}$. Then the corresponding empirical depth function $D_{n}$ almost surely uniformly to $D(\cdot,\nu)$ for $n$ going to infinity.
Proof 10
Due to the i.i.d. assumption and the law of large numbers, we know that for every $p\in\mathcal{P},\>\left\lVert\nu_{n}(p)-\nu(p)\right\rVert\overset{n\to\infty}{\to}0$ almost surely (a.s). Since #$\#\mathcal{P}$ is finite, we get that $\nu_{n}$ also converges a.s. uniformly to $\nu$. Finally, we use that $D_{n}$ and $D$ are both the same finite composition of $\nu_{n}$ and $\nu$, respectively, and we obtain $\sup_{p\in\mathcal{P}}\left\lVert D_{n}(p)-D(p)\right\rVert\overset{n\to\infty}{\to}0\text{ almost surely.}$.
@@ 6 Implementation
In this section, we discuss the difficulties and corresponding solution approaches in computing $D_{n}$. Therefore, let $\underline{p}=(p_{1},\ldots,p_{n})$ be a sample of posets. The naive approach is to just check all subsets of $\{p_{1},\ldots,p_{n}\}$ whether they are in $\mathscr{S}_{obs}:=\{S\in\mathscr{S}\mid\mbox{ all }p\in S\mbox{ are observed}\}$ is very time-consuming, especially since the subsets that are elements of $\mathscr{S}_{obs}$ can be very sparse in $2^{\{p_{1},\ldots,p_{n}\}}$. Moreover, it is difficult to test whether a subset is an element of $\mathscr{S}_{obs}$ or if it is not an element by taking Conditions (C1) and (C2) as a basis. Finally, even if we are able to compute the entire family $\mathscr{S}_{obs}$, computing all possible poset $p$ on a set $M$ to obtain some insights on $D_{n}$ is for even small $M$ with #$\#M\geq 8$ very computation intensive and for larger #$\#M$ it is currently not possible. [^16] Even computing the number of possible posets is not feasible for #$\#M>18$, see https://oeis.org/A001035 (Accessed 10.11.2023). We address all these issues in the next paragraphs.
@@
subject to $A_{poset}\>x\geq b_{poset}$
$A_{intersect}\>x\geq b_{intersect}$
$A_{union}\>x\geq b_{union}$
$x_{i}\in\{0,1\}\quad\text{for}\>i\in\mathscr{S}_{obs}\cup(M\times M)$
@@
where $\frac{w_{S}}{c_{n}}$ for $S\in\mathscr{S}_{obs}$ are the weights given by duplicates and the cardinality of $S$ together with the normalization constant, see Definition 4. The first #$\#\mathscr{S}_{obs}$ elements of $x$ describe the union-free generic set, and the next #${\#M\cdot\#M}$ elements guarantee that we only discuss posets. First, we need to ensure that only relations that define a poset are discussed. This is represented in Program (5) by the constraint $A_{poset}\>x\geq b_{poset}$. Therefore matrix $A_{poset}$ together with $b_{poset}$ represents only the reflexive, transitive, and antisymmetric constraints on $x_{(a,b)}$ with $(a,b)\in M\times M$. So in this part, the matrix entries of the first #$\#\mathscr{S}_{obs}$ columns are all zero. Second, we have to include in the binary linear program that a set $S\in\mathscr{S}_{obs}$ can only be included in the maximum if the considered poset lies in the intersection of all posets of $S$, see $A_{intersect}\>x\geq b_{intersect}$. Each row of this constraint corresponds to one ufg set $S\in\mathscr{S}$ and imposes a structure on $x_{(a,b)}$ with $(a,b)\in M\times M$ when this set $S$ is included in the maximization of the objective function. Similarly, we ensure with $\quad A_{union}\>x\geq b_{union}$ that a set $S\in\mathscr{S}_{obs}$ is only counted if the poset considered is a subset of the union of all posets in $S$. Now, computing the maximum leads to the highest depth value and a poset that has that highest depth value. However, if we are interested in the smallest depth value, this binary linear program does not work, because we have not yet enforced that a set $S\in\mathscr{S}_{obs}$ which has the currently analyzed poset in its closure must be represented in the calculation of the objective function. Therefore, we need to include further constraints so that when a poset is used to minimize the objective, every set $S\in\mathscr{S}_{obs}$ that has that poset in its closure is counted in the objective sum. The code with detailed comments can be found on GitHub, see Footnote 5.
@@
All in all, we improved the computation compared to the naive approach by using the knowledge provided in Section 5. By the two bounds on the cardinality of the sets in $\mathscr{S}_{obs}$ we get an idea of the worst and best case of the computation time. By further using Lemma 1 with Corollary 2 together the connectedness property, see Theorem 3 Part 3, we could decrease the computation time. Although we currently cannot calculate the exact amount of this reduction in general as this depends on the complexity of the data set used. Note that the upper bound using the VC dimension is not fixed, but depends on the structure of the data set. However, the bound given by Theorem 5 holds in general, but can be quite loose in some data situations. Finally, by stating a binary linear program, we achieved that even if not all posets are computable, we get the highest and lowest ufg depth values with representative posets. A summarizes the computation time and complexity of the two examples in the next section.